# RLAIF: Scaling Reinforcement Learning from Human Feedback with AI Feedback

### v2

@@ Abstract
Reinforcement learning from human feedback (RLHF) has proven effective in aligning large language models (LLMs) with human preferences. However, gathering high-quality human preference labels can be a time-consuming and expensive endeavor. RL from AI Feedback (RLAIF), introduced by Bai et al., offers a promising alternative that leverages a powerful off-the-shelf LLM to generate preferences in lieu of human annotators. Across the tasks of summarization, helpful dialogue generation, and harmless dialogue generation, RLAIF achieves comparable or superior performance to RLHF, as rated by human evaluators. Furthermore, RLAIF demonstrates the ability to outperform a supervised fine-tuned baseline even when the LLM preference labeler is the same size as the policy. In another experiment, directly prompting the LLM for reward scores achieves superior performance to the canonical RLAIF setup, where LLM preference labels are first distilled into a reward model. Finally, we conduct extensive studies on techniques for generating aligned AI preferences. Our results suggest that RLAIF can achieve human-level performance, offering a potential solution to the scalability limitations of RLHF.
@@
RLAIF: Scaling Reinforcement Learning from Human Feedback
with AI Feedback
Harrison Lee, Samrat Phatale, Hassan Mansoor, Thomas Mesnard, Johan Ferret, Kellie Lu, Colton Bishop, Ethan Hall, Victor Carbune, Abhinav Rastogi, Sushant Prakash Google Research {harrisonlee,samratph,hassan}@google.com
@@ 1 Introduction
Figure 1: Human evaluators strongly prefer RLAIF and RLHF over the SFT baseline for summarization and helpful dialogue generation. Their difference in win rates vs. SFT is not statistically significant. Furthermore, when compared head-to-head, RLAIF is equally preferred to RLHF. For harmless dialogue generation, RLAIF outperforms RLHF.
Reinforcement Learning from Human Feedback (RLHF) is an effective technique for aligning language models to human preferences (Stiennon et al., 2020; Ouyang et al., 2022). It is cited as one of the key drivers of success in modern conversational language models, such as ChatGPT (Liu et al., 2023) and Bard (Manyika, 2023). Training language models with reinforcement learning (RL) enables optimization on complex, sequence-level objectives that are not easily differentiable and therefore ill-suited for traditional supervised fine-tuning (SFT).
One obstacle for employing RLHF at scale is its dependence on high-quality human preference labels. This raises the question of whether artificially generated labels can be a viable substitute. Generating labels with large language models (LLMs) is one promising approach, as LLMs have shown a high degree of alignment with human judgment (Gilardi et al., 2023; Ding et al., 2023). Bai et al. (2022b) was the first effort to explore Reinforcement Learning from AI Feedback (RLAIF) [^1] This is distinct from “Constitutional AI”, which improves upon a supervised learning model through iteratively asking an LLM to generate better responses according to a a set of written value statements. Both were introduced in Bai et al. (2022b) and are sometimes conflated., where RL was conducted using a reward model trained on LLM preferences. Bai et al. (2022b) showed that utilizing a hybrid of human and AI preferences, in conjunction with their “Constitutional AI” self-revision technique, outperforms supervised fine-tuning for training a conversational assistant. However, it did not directly compare the efficacy of human vs. AI feedback, leaving the question of whether RLAIF can be a suitable alternative to RLHF unanswered.
@@
In this work, we study the impact of RLAIF and RLHF (see Figure 2) on three text generation tasks: summarization, helpful dialogue generation, and harmless dialogue generation. Our experiments show that RLAIF and RLHF are preferred by humans over the SFT baseline 71% and 73% of the time for summarization and 63% and 64% of the time for helpful dialogue generation, respectively, where the differences between RLAIF and RLHF win rates are not statistically significant. We also conduct a head-to-head comparison of RLAIF against RLHF and find that both policies are equally preferred [^2] The win rate for one policy vs. the other is not statistically significantly different from 50%. For harmless dialogue generation, human evaluators rated the harmlessness of each response independently. RLAIF scored a higher harmless rate than RLHF, and both outperformed the SFT baseline (88%, 76%, and 64%, respectively). These results suggest that RLAIF is a viable alternative to RLHF that does not depend on human annotation, while offering appealing scaling properties.
Additionally, we investigate two related questions. First, we explore whether RLAIF can improve upon a SFT policy when the LLM labeler has the same number of parameters as policy. Even in this scenario, RLAIF significantly improves over the SFT baseline. Second, we conduct an experiment where the off-the-shelf LLM is directly prompted for reward scores during RL, bypassing the step of distilling LLM preference labels into a reward model. This method achieves an even higher win rate over SFT than the canonical distillation method.
Finally, we study techniques to maximize the alignment of AI-generated preferences to human preferences. We find that soliciting chain-of-thought reasoning (Wei et al., 2022) consistently improves alignment, while using a detailed preamble and few-shot prompting (Brown et al., 2020) are only beneficial for certain tasks. We also conduct scaling experiments to examine the trade-off between the size of the LLM labeler and alignment with human preferences.
@@
    We demonstrate that RLAIF achieves comparable or superior performance to RLHF on the tasks of summarization, helpful dialogue generation, and harmless dialogue generation.
@@
    We show that RLAIF can improve upon a SFT policy even when the LLM labeler is the same size as the policy.
@@
    We find that directly prompting the LLM for reward scores during RL can outperform the canonical setup where a reward model is trained on LLM preferences.
@@
    We compare various techniques for generating AI labels and identify optimal settings for RLAIF practitioners.
@@ 2 Methodology
This section describes the techniques used to generate preferences with an LLM, how RL is conducted, and evaluation metrics. Preliminaries on RLHF are provided in Appendix A.
@@ 2.1 Preference Labeling with LLMs
We annotate preferences between pairs of candidates with an “off-the-shelf” LLM - a model pre-trained or instruction-tuned (Wei et al., 2021) for general usage but not fine-tuned for a specific downstream task. Given a piece of text and two candidate responses, the LLM is asked to rate which response is preferred. The prompt is structured as follows (examples in Tables 15 and 21):
@@
    Ending - Ending text to prompt the LLM (e.g. “Preferred Response=”)
@@
We experiment with two styles of preambles: “Base”, which essentially asks “which response is better?”, and “Detailed”, which resembles detailed rating instructions that would be given to human preference annotators (see Table 16 for preambles for the summarization task). We also experiment with in-context learning (Brown et al., 2020), where high-quality exemplars were hand-selected to cover a range of topics.
@@ 2.1.1 Addressing Position Bias
The order in which candidates are shown to an LLM can bias which candidate it prefers (Pezeshkpour and Hruschka, 2023; Wang et al., 2023). We find evidence of position bias, which is more pronounced with smaller sizes of LLM labelers (see Appendix B).
To mitigate position bias in preference labeling, we make two inferences for every pair of candidates, where the order in which candidates are presented to the LLM is reversed for the second inference. The results from both inferences are then averaged to obtain the final preference distribution.
2.1.2 Chain-of-thought Reasoning
Figure 3: An illustration of the process of obtaining AI-generated labels for summarization preferences. The LLM is first prompted to explain its thoughts on the quality of the two candidates (blue). The LLM’s response is then appended to the original prompt (orange) and fed to the LLM a second time to generate a preference distribution over “1” vs. “2” based on their log-probabilities (green).
We experiment with eliciting chain-of-thought (CoT) reasoning (Wei et al., 2022) from our AI labelers through a two-step inference procedure. First, we replace the Ending of the standard prompt (e.g. “Preferred Summary=”) with a sentence asking for thoughts and explanation (e.g. “Consider the coherence, accuracy, coverage, and overall quality of each summary and explain which one is better. Rationale:”) and then decode a response from the LLM. Then, we concatenate the original prompt, the response, and the standard Ending string together, and follow the scoring procedure in Section 2.1 to obtain a preference distribution. See Figure 3 for an illustration.
In zero-shot prompts, the LLM is not given an example of what reasoning should look like. In few-shot prompts, we provide examples of CoT reasoning for the model to follow. See Tables 17 and 18 for examples.
@@ 2.2 Reinforcement Learning from AI Feedback
2.2.1 Distilled RLAIF
We describe our adaptation of the canonical RLAIF setup below, which we also refer to as “distilled RLAIF”. Unless otherwise mentioned, RLAIF is carried out using this method.
After labeling preferences with an LLM, a reward model (RM) is trained on these labels. Since our approach produces soft labels (e.g. $[0.6,0.4]$), we apply a cross-entropy loss to the softmax of the reward scores generated by the RM. The softmax converts the RM scores into a probability distribution. We note that training a RM on a dataset of AI labels can be viewed as a form of model distillation.
Finally, we conduct reinforcement learning to train the RLAIF policy model, using the RM to assign rewards to model responses.
2.2.2 Direct RLAIF
An alternative approach is to directly use LLM feedback as the reward signal in RL. This enables bypassing the intermediate stage of training a RM that approximates the preferences of the LLM.
The LLM is prompted to rate the quality of a generation between 1 and 10. Similar to the format mentioned in Section 2.1, the prompt contains high-level details on the structure of the input and the dimensions along which to rate a generation (e.g. factuality, coherence). Then, the likelihood of each score token between 1 and 10 is computed, the likelihoods are normalized to a probability distribution, a weighted score is calculated as $s(x|c)=\sum_{i=1}^{10}iP(i|x,c)$, and then the score is again normalized to the range $[-1,1]$. Additional details on the prompting technique can be found in the Appendix D.
Finally, RL is conduct RL in a similar manner to “distilled RLAIF”, where the direct score is used as reward instead of the score from a RM. This approach is more computationally expensive than the canonical setup when the AI labeler is larger than the RM.
@@ 2.3 Evaluation
AI Labeler Alignment measures the accuracy of AI-labeled preferences with respect to human preferences. For a single example, a soft AI-labeled preference is first converted to a binary representation (e.g. $[0.6,0.4]\rightarrow[1,0]$). Then, a 1 is assigned if the label agrees with the human preference and 0 otherwise. The alignment accuracy $z_{acc}$ can be expressed as follows:
@@
where $D$ is the size of the preference dataset, $P^{AI}\in\mathbb{R}^{D\times 2}$ is the matrix of soft AI preferences, and $p^{human}\in\mathbb{R}^{D}$ is the corresponding vector of human preferences, containing elements 00 or $1$ to denote whether the first or second response is preferred, respectively.
Win Rate evaluates the end-to-end quality of two policies by measuring how often one policy is preferred by human annotators over another. Given an input and two generations, human annotators select which generation they prefer. The percentage of instances where policy $A$ is preferred over policy $B$ is referred to as the “win rate of A vs. B”. A 50% win rate indicates that $A$ and $B$ are equally preferred.
@@ 3.1 Datasets
  • Reddit TL;DR (Stiennon et al., 2020) - posts from Reddit333www.reddit.com accompanied by summaries of the posts.
@@
  • Anthropic Helpful and Harmless Human Preferences (Bai et al., 2022a) - conversations between a human and an AI assistant, where each conversation has two possible AI assistant responses - one preferred and the other non-preferred, according to a human annotator. Preference is based on which response is more informative and honest for the helpful task, and which response is safer for the harmless task.
@@
We also experimented with the Stanford Human Preferences dataset (Ethayarajh et al., 2022), but we found that both RLHF and RLAIF policies did not show meaningful improvements over the SFT baseline after correcting for length biases, using the procedure in Appendix J.
@@ 3.2 LLM Labeling
To enable fast experiment iteration when evaluating AI labeling techniques, we randomly downsampled the training split of each preference dataset. For summarization, an additional filter was applied to only include examples where human annotators preferred one summary over the other with high confidence [^4] This follows the evaluation procedure in Stiennon et al. (2020). Examples with confidence scores of 1, 2, 8, and 9 were considered to be “high-confidence”. After downsampling and filtering, there were roughly 3-4k examples for each task [^5] We sample 15%, 10%, and 10% of the training splits for summarization, helpful dialogue generation, and harmless dialogue generation, respectively.. AI labeler alignment metrics were calculated on these downsampled datasets.
PaLM 2 (Google et al., 2023) is used as the LLM for labeling preferences. The versions used are instruction-tuned but not previously trained with RL. Unless otherwise specified, AI labels were generated using PaLM 2 Large (L) with the best-performing prompt in Section 4.4. For more details on LLM labeling, see Appendix D.
@@ 3.3 Model Training
RMs are also derived from PaLM 2 XS. RMs are fine-tuned on the entire training split of the corresponding preference dataset, where the label is the AI preference for AI feedback RMs and the original human preference label in the dataset for human feedback RMs. RM accuracies can be found in Appendix G.
In the RL phase, the policy is trained with a modified version of REINFORCE (Williams, 1992) adapted to the language modeling domain. While many recent works use Proximal Policy Optimization (PPO) (Schulman et al., 2017) - a related method that adds a few techniques to make training more conservative and stable (e.g. clipping the objective function), we use REINFORCE with a baseline given that it is simpler yet still effective for the problem at hand. Both policy and value models are initialized from the SFT model. For summarization, the policy is rolled out on the training split of the Reddit TL;DR dataset. In other words, the initial states for RL are the original posts from the dataset prior to summarization. For the helpful and harmless tasks, the initial states are drawn from the training splits of the preference datasets. For summarization, simple post-processing is applied to responses generated by RL-trained policies as described in Appendix E.
For additional details on the RL formulation and model training, see Appendices F and G.

### v3

# RLAIF vs. RLHF: Scaling Reinforcement Learning from Human Feedback with AI Feedback

@@ Abstract
Reinforcement learning from human feedback (RLHF) has proven effective in aligning large language models (LLMs) with human preferences, but gathering high-quality preference labels is expensive. RL from AI Feedback (RLAIF), introduced in Bai et al. (2022b), offers a promising alternative that trains the reward model (RM) on preferences generated by an off-the-shelf LLM. Across the tasks of summarization, helpful dialogue generation, and harmless dialogue generation, we show that RLAIF achieves comparable performance to RLHF. Furthermore, we take a step towards “self-improvement” by demonstrating that RLAIF can outperform a supervised fine-tuned baseline even when the AI labeler is the same size as the policy, or even the exact same checkpoint as the initial policy. Finally, we introduce direct-RLAIF (d-RLAIF) - a technique that circumvents RM training by obtaining rewards directly from an off-the-shelf LLM during RL, which achieves superior performance to canonical RLAIF. Our results suggest that RLAIF can achieve performance on-par with using human feedback, offering a potential solution to the scalability limitations of RLHF.
Reinforcement Learning, Large Language Models, Alignment with Human Feedback, Alignment from AI Feedback, Self-improvement
@@
Figure 1: Human evaluators strongly prefer RLAIF and RLHF over the SFT baseline for summarization and helpful dialogue generation. Furthermore, when compared head-to-head, RLAIF is equally preferred to RLHF. For harmless dialogue generation, RLAIF outperforms RLHF.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) is an effective technique for aligning language models to human preferences (Stiennon et al., 2020; Ouyang et al., 2022). It is cited as one of the key drivers of success in modern conversational language models, such as ChatGPT (Liu et al., 2023) and Bard (Manyika, 2023). A key advantage of training language models with reinforcement learning (RL) is that it enables optimization on complex, sequence-level objectives that are not easily differentiable and therefore ill-suited for traditional supervised fine-tuning (SFT).
@@
One obstacle for employing RLHF at scale is its dependence on high-quality human preference labels. Modern large language models (LLMs) have shown a high degree of alignment with human judgment (Gilardi et al., 2023; Ding et al., 2023), suggesting that LLM-generated preference labels may be a viable substitute for human labels. Bai et al. (2022b) was the first effort to explore Reinforcement Learning from AI Feedback (RLAIF), where RL was conducted using a reward model trained on a hybrid of human and AI preferences. In conjunction with their “Constitutional AI” self-revision technique, their final policy outperformed supervised fine-tuning for training a conversational assistant. However, it did not directly compare the efficacy of human vs. AI feedback, leaving the question of whether RLAIF can be a suitable alternative to RLHF unanswered.
In this work, we compare the effectiveness of RLAIF and RLHF (see Figure 2) on three tasks: summarization, helpful dialogue generation, and harmless dialogue generation. Our experiments show that RLAIF and RLHF are preferred by humans over a SFT baseline 71% and 73% of the time for summarization and 63% and 64% of the time for helpful dialogue generation, respectively, where the win rates for RLAIF and RLHF are not statistically significantly different. Furthermore, in a head-to-head comparison of RLAIF against RLHF, both policies are equally preferred [^1] The win rate for one policy over the other is not statistically significantly different from 50%. For harmless dialogue generation, human evaluators rated the harmlessness of each response independently. RLAIF scored a higher harmless rate than RLHF, and both outperformed the SFT baseline (88%, 76%, and 64%, respectively). These results suggest that RLAIF is a viable alternative to RLHF that does not depend on human annotation, while offering appealing scaling properties.
Additionally, we conduct two related studies. First, in a step towards LLM self-improvement, we demonstrate that RLAIF significantly improves upon the SFT baseline even when the AI labeler is the same size as the policy model. Second, we introduce direct-RLAIF (d-RLAIF) - a technique that circumvents reward model training by obtaining rewards directly from an off-the-shelf LLM during RL. In our experiments, d-RLAIF matches or outperforms canonical RLAIF. Furthermore, for the task of helpful dialogue generation, the initial policy and the LLM providing rewards are the same model checkpoint, demonstrating an instance of strict LLM self-improvement.
Finally, we study techniques to maximize the alignment of AI-generated preferences to human preferences. We find that soliciting chain-of-thought reasoning (Wei et al., 2022) consistently improves alignment, while using a detailed preamble and few-shot prompting (Brown et al., 2020) are only beneficial for certain tasks. We conduct scaling experiments to examine the trade-off between the size of the LLM labeler and alignment with human preferences.
@@
    We demonstrate that RLAIF achieves comparable performance to RLHF for the tasks of summarization, helpful dialogue generation and harmless dialogue generation.
@@
    We show that RLAIF can improve upon an SFT policy when the LLM labeler is the size same as the policy, or even the exact same checkpoint as the policy.
@@
    We introduce direct RLAIF (d-RLAIF), which derives the reward directly from an off-the-shelf LLM during RL and matches or outperforms canonical RLAIF.
@@
    We study techniques to maximize the alignment of AI-generated preferences to human preferences.
Figure 3: An illustration of the process to obtain AI-generated preference labels for summarization. The LLM is first prompted to explain its thoughts on the quality of the two candidates (blue). The response (orange) is then appended to the first prompt, and together they form the second prompt used to generate a preference distribution over “1” vs. “2” (green).
@@ 2 Methodology
This section describes the techniques used to generate preferences with an LLM, the reinforcement learning setups, and evaluation metrics. RLHF preliminaries are provided in Appendix A.
@@ 2.1 Preference Labeling with LLMs
We annotate preferences with an “off-the-shelf” LLM - a model pre-trained or instruction-tuned (Wei et al., 2021) for general usage but not fine-tuned for a specific downstream task. Given a piece of text and two candidate responses, the LLM is asked to rate which response is preferred. The prompt is structured as follows (examples in Tables 15 and 21):
@@
    Ending - The ending text to prompt the LLM (e.g. “Preferred Response=”)
@@
We experiment with two styles of preambles: “Base”, which essentially asks which response is better, and “Detailed”, which resembles detailed rating instructions typically given to human annotators (see Table 16 for preambles used in the summarization task). We also experiment with in-context learning (Brown et al., 2020), using high-quality exemplars hand-selected to cover a range of topics.
@@ 2.1.1 Addressing Position Bias
The order in which candidates are shown to an LLM can bias which candidate it prefers (Pezeshkpour & Hruschka, 2023; Wang et al., 2023). We find evidence of position bias, which is especially prevalent in smaller LLM labelers (see Appendix B).
To mitigate the effect of position bias, two inferences are made for every pair of candidates, where the order in which candidates are presented to the LLM is reversed for the second inference. The results from both inferences are then averaged to obtain the final preference distribution.
2.1.2 Eliciting Chain-of-thought Reasoning
We experiment with eliciting chain-of-thought (CoT) reasoning (Wei et al., 2022) from our AI labelers through a two-step inference procedure. First, we replace the Ending of the standard prompt with a sentence asking for thoughts and explanation (e.g. “Consider the coherence, accuracy, coverage, and overall quality of each summary and explain which one is better. Rationale:”) and decode a response from the LLM. Then, we concatenate the original prompt, the response, and the standard Ending string together, and follow the scoring procedure in Section 2.1 to obtain a preference distribution. Figure 3 illustrates this process.
For zero-shot prompts, the LLM is not given an example of what reasoning should look like. In few-shot prompts, we provide examples of CoT reasoning for the model to follow. See Tables 17 and 18 for examples.
Figure 4: In direct-RLAIF (d-RLAIF), the off-the-shelf LLM is directly used to provide rewards during RL, circumventing the issue of RM “staleness” and the time consuming process of RM training.
@@ 2.2 Reinforcement Learning from AI Feedback
2.2.1 Canonical RLAIF
We describe our adaptation of the canonical RLAIF setup below. Unless otherwise mentioned, RLAIF is carried out using this method.
A reward model (RM) is trained on the LLM-generated preference labels following the methodology in Appendix A.2. Since our approach produces soft labels (e.g. $[0.6,0.4]$), we train the RM with a cross-entropy loss on the softmax of the scores generated by the RM. The softmax converts the RM scores into a probability distribution. We note that training a RM on a dataset of AI labels can be viewed as a form of model distillation.
Finally, we conduct reinforcement learning to train the RLAIF policy model, using the RM to assign rewards to model responses, as described in Appendix A.3.
2.2.2 Direct-RLAIF (d-RLAIF)
One issue with RLAIF is that the reward model may become “stale” as the policy is trained. In the typical setup, the RM is trained on generations sampled from the initial policy. As the policy is trained, the generated trajectories become increasingly out-of-distribution from the dataset the RM was trained on, leading to suboptimal performance (Bai et al., 2022a). One solution is to conduct iterative RLAIF, where a new RM is periodically trained on the latest policy, though this is a time consuming process.
We introduce direct-RLAIF (d-RLAIF) - a simple alternative to canonical RLAIF that directly uses LLM feedback as the reward signal in RL. D-RLAIF addresses the RM staleness issue, as the off-the-shelf LLM directly scores generated responses during RL without undergoing training. Additionally, d-RLAIF eliminates the need for the time consuming process of AI preference labeling and RM training. Figure 4 depicts this process.
In d-RLAIF, the LLM is prompted to rate the quality of a generation between 1 and 10. Similar to Section 2.1, the prompt instructs the LLM on how to rate a generation. Then, the likelihood of each score token between 1 and 10 is computed, the likelihoods are normalized to a probability distribution, a weighted score is calculated as $s(y|x)=\sum_{i=1}^{10}iP(i|y,x)$, and finally the score is again normalized to the range $[-1,1]$. Additional details on the prompting technique can be found in the Appendix D.
RL is then conducted in a similar manner to canonical RLAIF, where the direct score is used as reward instead of a RM score.
@@ 2.3 Evaluation
AI Labeler Alignment measures the accuracy of AI-labeled preferences with respect to human preferences. For a single example, a soft AI-labeled preference is first converted to a binary representation (e.g. $[0.6,0.4]\rightarrow[1,0]$). Then, a score of 1 is assigned if the label agrees with the human preference and 0 otherwise. The alignment accuracy $z_{acc}$ can be expressed as follows:
@@
where $D$ is the size of the preference dataset, $P^{AI}\in\mathbb{R}^{D\times 2}$ is the matrix of soft AI preferences, and $p^{H}\in\mathbb{R}^{D}$ is the corresponding vector of human preferences, containing elements 00 or $1$ to denote whether the first or second response is preferred, respectively.
Win Rate evaluates the end-to-end quality of two policies by measuring how often one policy is preferred by human annotators over another. Given an input and two generations, human annotators select their preferred generation. The percentage of instances where policy $A$ is preferred over policy $B$ is referred to as the “win rate of A vs. B”. A 50% win rate indicates that $A$ and $B$ are equally preferred.
@@ 3.1 Datasets
  • Reddit TL;DR (Stiennon et al., 2020) - posts from Reddit222www.reddit.com accompanied by summaries of the posts.
@@
  • Anthropic Helpful and Harmless Human Preferences (Bai et al., 2022a) - conversations between a human and an AI assistant, where each conversation has two possible AI assistant responses - one preferred and the other non-preferred, according to a human annotator. Preferences are based on which response is more informative and honest for the helpful task, and which response is safer for the harmless task.
@@
We also explored the Stanford Human Preferences dataset (Ethayarajh et al., 2022), but we found that both RLHF and RLAIF policies did not show meaningful improvements over the SFT baseline after correcting for length biases as described in Appendix J.
@@ 3.2 LLM Labeling
To enable fast experiment iteration when evaluating AI labeling techniques, we randomly downsampled the training split of each preference dataset. For summarization, an additional filter was applied to only include examples where human annotators preferred one summary over the other with high confidence [^3] This follows the evaluation procedure in Stiennon et al. (2020). Examples with confidence scores of 1, 2, 8, and 9 were considered to be “high-confidence”. After downsampling and filtering, there remained 3-4k examples for each task [^4] We sampled 15%, 10%, and 10% of the training splits for summarization, helpful dialogue generation, and harmless dialogue generation, respectively.. AI labeler alignment was calculated on these downsampled datasets.
We use the PaLM 2 (Google et al., 2023) family of models for labeling preferences. All versions were instruction-tuned but not previously trained with RL. Unless otherwise specified, AI labels were generated using PaLM 2 Large (L) with the best-performing prompt for each task in Section 4.4. For more details on LLM labeling, see Appendix D.
@@ 3.3 Model Training
All RMs are trained from PaLM 2 XS checkpoints. Each RM is fine-tuned on the full training split of a preference dataset, where the label is the AI preference for AI feedback RMs and the original human preference label for human feedback RMs. RM accuracies can be found in Appendix G.
In the RL phase, the policy is trained with a modified version of REINFORCE (Williams, 1992) adapted to the language modeling domain (see Appendix E). While many recent works use Proximal Policy Optimization (PPO) (Schulman et al., 2017), we use REINFORCE with a baseline given that it is simpler yet still effective for the problem at hand. Both policy and value models are initialized from the SFT model. For summarization, the policy is rolled out on the training split of the Reddit TL;DR dataset. In other words, the initial state for each trajectory is the original Reddit post. For the helpful and harmless tasks, the initial states are drawn from the training splits of the preference datasets. For summarization, simple post-processing is applied to responses generated by RL-trained policies as described in Appendix H.
For additional details on model training, see Appendix F.